Using traditional methods it takes 11.3 hours to receive a basic flying license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique on 20 students and observed that they had a mean of 11.6 hours with a standard deviation of 1.1. Is there evidence at the 0.05 level that the technique performs differently than the traditional method? Assume the population distribution is approximately normal.

Step 1 of 5: State the null and alternative hypotheses.

Step 2 of 5:

Find the value of the test statistic. Round your answer to three decimal places.

Step 3 of 5:

Specify if the test is one-tailed or two-tailed.

One-Tailed Test

OR

Two-Tailed Test

Step 4 of 5:

Determine the decision rule for rejecting the null hypothesis. Round your answer to three decimal places.

Reject H0 if |t|>  

Step 5 of 5:

Make the decision to reject or fail to reject the null hypothesis.

Reject Null Hypothesis

OR

Fail to Reject Null Hypothesis
Answer #1

This is the two tailed test .

The null and alternative hypothesis is

H0 : \mu = 11.3

Ha : \mu \neq 11.3

Test statistic = t

= (\bar x - \mu ) / s / \sqrt n

= (11.6 - 11.3) / 1.1 / \sqrt 20

Test statistic = 1.220

Two tailed test

df = 20 - 1 = 19

Reject H0 if |t| > 2.093

Test statistics < critical value

Fail to reject the null hypothesis .